# Planning-Driven Programming: A Large Language Model Programming Workflow

### v1

@@ A.3 Additional Ablation Study
Table 8: The percentage of problems where Self-Debugging (+Expl) (SD), LDB, LPW, and SLPW generated programs solve the visible tests but fail the hidden tests, out of total failed problems for each method on HumanEval and MBPP, with GPT-3.5 as the backbone.
@@
Table 9: The percentage of problems where Self-Debugging (+Expl) (SD), LDB, LPW, and SLPW generated programs pass the visible tests but fail the hidden tests, out of a total of 164 problems in HumanEval and 500 problems in MBPP, with GPT-3.5 as the backbone.
@@
A.4 Analysis of Unsolved Problems for GPT-3.5
LPW and SLPW achieve state-of-the-art performance among evaluated methods and show dominance compared to other LLM debuggers. We categorize the failure reasons for LPW and SLPW on HumanEval with GPT-3.5 into 5 types. Table 7 compares the percentage of different failure reasons out of the total unsolved problems for LPW and SLPW based on authors’ manual review. Approximately half of the errors result from the No Code type, where the generated solution plan fails to be verified on the visible tests, or the resulting verification includes incorrect intermediate outputs in the solution generation phase, leading to failure after reaching the maximum iteration threshold. The second most common reason is Missing Conditions, originating from the same issues in the plan and plan verification. Notably, LPW generated program solutions contain no logic errors, whereas SLPW produces only one program with a logic error. Both SLPW and LPW fail in the 91st problem, where the generated programs are unable to solve the hidden tests due to deviations from the plan and plan verification (Differ from Intended Solution). The plan verification clearly specifies splitting the input string into sentences using delimiters “.”, “?” or “!”, but the generated code only handles the full stop case and ignores “?” and “!”.
@@
Tables 9 and 9 show the percentage of problems where Self-Debugging (+Expl) (SD), LDB, LPW, and SLPW generated program solutions pass the visible tests but fail the hidden tests, out of respectively failed problems and the total number of problems in the HumanEval and MBPP benchmarks using GPT-3.5 as the backbone. In Table 9, more than 40% of failures in LPW and SLPW result from solving the visible tests only on the HumanEval benchmark, since except for the No Code category, other reasons discussed in Table 7 could contribute to this issue. In contrast, less than 30% of problems in SD and LDB experience this issue on HumanEval as the larger number of failed problems in these two methods. In Table 9, all evaluated approaches show similar percentages on the MBPP benchmark, with the remaining failures arising from different reasons. We note that all methods tend to address visible tests only on the same set of problems in both the HumanEval and MBPP benchmarks, resulting in the similar percentage in each benchmark out of the total number of problems, as shown in Table 9. Meanwhile, all methods are prone to addressing visible tests only on MBPP rather than on HumanEval as indicated in Table 9. Compared to the detailed problem descriptions in HumanEval, the problem descriptions in MBPP are concise but lack clarity. For example, Figure 6 illustrates two problems in MBPP where LPW and SLPW generated solutions are tailored to the visible tests but deviate significantly from the canonical solution.
A.5 Accuracy of solution plans and plan verifications using GPT-3.5
We manually investigate the accuracy of solution plans and plan verifications generated by GPT-3.5 on the HumanEval benchmark, and the results are presented in Table 10. Overall, GPT-3.5 generates the correct solution plans and plan verifications in natural language for majority of problems. In LPW and SLPW, GPT-3.5 successfully produces plans and plan verifications for more than 94% of the problems. GPT-3.5 generates the correct plans for around 93% of the problems and achieves the similar accuracy for plan verifications. One common issue in the LLM-generated plan is the omission of certain conditions. For example, solution plan frequently overlooks uppercase situations and negative numbers. We note that the LLM-generated plan verification closely adheres to the solution plan. When the plan is accurate, the verification process strictly follows the plan logic, resulting in a correct analysis. Conversely, if the plan contains logical errors or omits edge cases, the verification process replicates these mistakes.
We further manually explore the relationship between plan verification and program solution on the HumanEval benchmark with GPT-3.5. Table 11 evaluates the conditional probabilities between wrong code and wrong plan verification, as well as between correct code and correct plan verification. Typically, in LPW and SLPW, accurate plan verification significantly contributes to an accurate program solution, and vice versa. In LPW and SLPW, GPT-3.5 generates program solutions based on plans and plan verifications. Therefore, any accurate descriptions or mistakes, including missed conditions, in the plan and plan verification are propagated to the code. For SLPW, one problem’s verification is classified as wrong, while the subsequently generated program is correct, resulting in the value in the first column being less than 100%. This discrepancy arises since the plan verification contains correct logic but also includes numerous noisy messages, leading to its manual classification as wrong. Despite this, the generated program is correct due to the sound underlying logic in the plan verification. When plan verifications are accurate, over 95% of program solutions are correct in LPW and SLPW. The remaining incorrect code instances result from unclear condition statements for hidden tests in plan verification, leading to an error program solution.
The results from Tables 10 and 11 highlight the impressive capabilities of LLMs in tackling text-to-code generation tasks when outputs are represented in natural language. Plan and plan verification generation accuracy is typically higher than code generation accuracy, underscoring the rationale behind LPW and SLPW, which produce the high-quality program solution by leveraging plan and plan verification. It is worth exploring methods to help LLMs overcome the challenges of translating natural language solutions into programs, given the strict lexical, grammatical, and semantic constraints. Additionally, exploring other types of natural language solution representations could improve code generation accuracy.
@@
Table 10: Percentage of problems where the LLM successfully generates the valid plans and plan verifications in the solution generation phase (first column); percentage of problems where the LLM-generated plans are manually classified as correct (middle column), considering no plan cases; and percentage of problems where the LLM-generated plan verifications are manually classified as correct (last column), considering no plan verification cases. SLPW generates multiple plans and plan verifications for each problem, and we consider them correct only when all are classified as correct. All percentages are reported using GPT-3.5 as the backbone on the HumanEval benchmark, with a total of 164 problems.
@@
Table 11: The relationship between LLM-generated code solutions and plan verifications on the HumanEval benchmark with GPT-3.5. The first column shows the percentage of problems where the LLM generates incorrect code solutions when plan verifications are incorrect; the second column shows the percentage of problems where correct code solutions are generated when plan verifications are correct. For SLPW, only the plan verification used to generate the final output code is manually evaluated.
A.6 Refinement Consistency in LPW and SLPW
@@
LPW and SLPW allow multiple rounds of debugging to refine code based on error analysis generated by comparing the code execution trace and plan verification on the failed visible test. Additionally, LPW and SLPW query LLMs to generate refined code accompanied by a refinement explanation, detailing the modifications implemented to address the errors identified in the error analysis. For instance, Figures 9 and 5 illustrate two HumanEval problems where LPW successfully generates the correct program through refinements informed by the error analysis using the GPT-3.5 backbone. We note that in LPW and SLPW, if the refined code is irrelevant to the error analysis, the entire debugging process degrades to a simple sampling approach, contradicting our original intent. As a result, we manually evaluate the debugging consistency among the generated error analysis (part (e)), the refined code (part (f)), and the refinement explanation (part (g)), as exampled in Figure 9. In LPW, only one refined code deviates from the error analysis yet still produces the correct solution, across all problems solved through debugging. SLPW achieves perfect consistency between the error analysis and the refined code. These results validate the effectiveness of the debugging steps in the code implementation phase for both LPW and SLPW, where the meaningful error analysis enables LLMs to produce the correct program with precise refinements.
A.7 Analysis of Unsolved Problems for GPT-4o
A.7.1 HumanEval
@@
SLPW achieves 98.2% Pass@1 accuracy on HumanEval with the GPT-4o backbone, indicating only 3 unsolvable problems. We further investigate the reasons behind GPT-4o’s failures on the 91st, 132nd, and 145th problems as shown in Figures 10, 11, and 12, and attempt to generate the correct program solutions. The 91st problem fails since GPT-4o misinterprets the linguistic distinction between the word and the letter; the 132nd problem’s ambiguous description challenges GPT-4o; and the incomplete description of the 145th problem leads to failed plan verifications. GPT-4o successfully generates correct program solutions for 2 out of 3 problems, achieving 99.4% Pass@1 accuracy, by involving an additional visible test to validate the intended solution for the 91st problem and providing a comprehensive problem description for the 145th problem.
@@
Figure 12 illustrates the 132nd problem, where ambiguity in the problem description (part (a)) challenges GPT-4o. The problem description lacks clarity on “a valid subsequence of brackets” and fails to specify the meaning of “one bracket in the subsequence is nested”. We deduce the intended problem description by prompting GPT-4o with a canonical solution (part (d)). However, the LLM-generated description remains unclear and results in various erroneous programs. Furthermore, adding typically failed hidden tests to the visible tests (part (b)) is also ineffective in clarifying the correct logic. We acknowledge that a clearer description might contribute to the correct program. However, some problems are inherently difficult to describe accurately in natural language without careful organization, posing challenges for LLMs.
@@
Table 12: The percentage of problems where LDB, LPW, and SLPW generated programs solve the visible tests but fail the hidden tests, out of total failed problems for each method in APPS and CodeContests, with GPT-4o as the backbone.
@@
Table 13: The percentage of problems where LDB, LPW, and SLPW generated programs pass the visible tests but fail the hidden tests, out of a total of 139 problems in APPS and 150 problems in CodeContests, with GPT-4o as the backbone.
@@
Figure 7: Pass@1 accuracy of Baseline, LDB, LPW and SLPW across different difficulty levels, Introductory, Interview, and Competition, on the APPS benchmark when using GPT-4o as the LLM backbone.
Figure 8: Pass@1 accuracy vs. average token cost per program for LDB, LPW, and SLPW on the APPS benchmark using GPT-4o as the LLM backbone.
A.7.2 APPS and CodeContests
@@
We preprocess problem descriptions in APPS and CodeContests, as shown in Figures 13 and 14, to maintain consistency in the input data structure. LPW and SLPW demonstrate significant improvements on APPS and CodeContests, exceeding around 10% and 5% Pass@1 accuracy, respectively, compared to LDB with GPT-4o. However, in contrast to their performance on the HumanEval and MBPP benchmarks, where LPW and SLPW achieve over 97% and 84% Pass@1 accuracy, the 62% accuracy on APPS and 34% accuracy on CodeContests indicate that even for the advanced LLM GPT-4o, code generation remains challenging when addressing complicated programming problems, such as those encountered in collegiate programming competitions like IOI and ACM (Hendrycks et al., 2021).
@@
Figure 8 compares the Pass@1 accuracy of LDB, LPW, and SLPW across different difficulty levels, Introductory, Interview, and Competition, on the APPS benchmark using GPT-4o. LPW and SLPW consistently dominate in Pass@1 accuracy across all difficulty levels. LPW and SLPW show strong performance on the relatively easier levels, i.e., Introductory and Interview, surpassing LDB by around 9% and 15% accuracy, respectively, and outperforming Baseline by over 20% accuracy. For the problems belonging to the most challenging level, Competition, LPW and SLPW achieve 34.8% accuracy, compared to 28.3% for LDB and 17.4% for Baseline. However, all approaches experience a substantial decrease at the Competition level, emphasizing the need for further improvements. Figure 8 compares the Pass@1 accuracy and average token cost per program for LDB, LPW, and SLPW on APPS using GPT-4o. The results align with expectations: LDB uses the fewest tokens, as it only refines seed programs that fail visible tests, generated through 0-shot prompting; LPW achieves better accuracy but requires additional tokens for generating the plan and plan verification; and SLPW achieves the highest accuracy but consumes the most tokens due to the creation of multiple plans, plan verifications and programs.

### v2

@@ D.1 Performance Across Different Difficulty Levels
Table 7: The percentage of problems where MapCoder, Self-Debugging (+Expl) (SD), LDB, and LPW generated programs solve the visible tests but fail the hidden tests, out of total failed problems for each method on HumanEval and MBPP, with GPT-3.5 as the backbone.
@@
Table 8: The percentage of problems where MapCoder, Self-Debugging (+Expl) (SD), LDB, and LPW generated programs pass the visible tests but fail the hidden tests, out of a total of 164 problems in HumanEval and 500 problems in MBPP, with GPT-3.5 as the backbone.
D.2 Failure Reasons
LPW achieves state-of-the-art performance among the evaluated methods and demonstrates superiority over other LLM debuggers. We categorize the failure reasons for LPW on HumanEval with GPT-3.5 into 4 types. Table 6 compares the percentage of different failure reasons out of the total unsolved problems for LPW based on authors’ manual review. In LPW, half of the errors result from the No Code type, where the generated solution plan fails to be verified on the visible tests, or the resulting verification includes incorrect intermediate outputs in the solution generation phase, leading to failure after reaching the maximum iteration threshold. The second most common reason is Missing Conditions, originating from the same issues in the plan and plan verification. For LPW, 5.6% of failures result from the generated program solution differing from the plan and plan verification (Differ from Intended Solution). For example, LPW fails in the 91st problem, where the generated program is unable to solve the hidden tests due to deviations from the plan and plan verification. The plan verification clearly specifies splitting the input string into sentences using delimiters “.”, “?” or “!”, but the generated code only handles the full stop case and ignores “?” and “!”.
@@ D.2 Failure Reasons
D.3 Failure on Hidden Tests
@@
Tables 7 and 8 show the percentage of problems where MapCoder, Self-Debugging (+Expl) (SD), LDB, and LPW generated program solutions pass the visible tests but fail the hidden tests, out of respectively failed problems and the total number of problems in the HumanEval and MBPP benchmarks using GPT-3.5 as the backbone. In Table 7, 44.4% of failures in LPW result from solving the visible tests only on the HumanEval benchmark, since except for the No Code category, other reasons discussed in Table 6 could contribute to this issue. In contrast, less than 30% of failures in MapCoder, SD and LDB experience this issue on HumanEval as the larger number of failed problems in these methods. In Table 7, MapCoder demonstrates a higher likelihood of passing only the visible tests, while all other evaluated approaches show similar percentages on the MBPP benchmark, with the remaining failures arising from different reasons. We note that all methods tend to address visible tests only on the same set of problems in the HumanEval benchmark, resulting in the similar percentage out of the total number of problems, as shown in Table 8. In contrast, on the MBPP benchmark, MapCoder exhibits the highest rate of passing only the visible tests out of the total number of problems, consistent with the result in Table 7, while LPW demonstrates the lowest rate, as shown in Table 8. Meanwhile, all methods are prone to addressing visible tests only on MBPP rather than on HumanEval as indicated in Table 8. Compared to the detailed problem descriptions in HumanEval, the problem descriptions in MBPP are concise but lack clarity. For example, Figure 11 illustrates two problems in MBPP where LPW generated solutions are tailored to the visible tests but deviate significantly from the canonical solution.
@@
Table 9: Percentage of problems where the LLM successfully generates the valid plans and plan verifications in the solution generation phase (first column); percentage of problems where the LLM-generated plans are manually classified as correct (middle column), considering no plan cases; and percentage of problems where the LLM-generated plan verifications are manually classified as correct (last column), considering no plan verification cases. All percentages are reported using GPT-3.5 as the backbone on the HumanEval benchmark, with a total of 164 problems.
@@
Table 10: The relationship between LLM-generated code solutions and plan verifications on the HumanEval benchmark with GPT-3.5. The first column shows the percentage of problems where the LLM generates incorrect code solutions when plan verifications are incorrect; the second column shows the percentage of problems where correct code solutions are generated when plan verifications are correct.
Appendix E Accuracy of Plans, Plan Verifications, and Programs in LPW using GPT-3.5
E.1 Plans and Plan Verifications
We manually investigate the accuracy of solution plans and plan verifications generated by GPT-3.5 on the HumanEval benchmark, and the results are presented in Table 9. Overall, GPT-3.5 generates the correct solution plans and plan verifications in natural language for majority of problems. In LPW, GPT-3.5 successfully produces plans and plan verifications for 94.5% of the problems. GPT-3.5 generates the correct plans for 92.7% of the problems and achieves the same accuracy for plan verifications. A common issue in the LLM-generated plan is the omission of certain conditions. For example, solution plan frequently overlooks uppercase situations and negative numbers. We note that the LLM-generated plan verification closely adheres to the solution plan. When the plan is accurate, the verification process strictly follows the plan logic, resulting in a correct analysis. Conversely, if the plan contains logical errors or omits edge cases, the verification process replicates these mistakes. Specifically, for LPW, all correct plans lead to accurate plan verifications, and vice versa.
E.2 Plan Verifications and Programs
We further manually explore the relationship between plan verification and program solution on the HumanEval benchmark with GPT-3.5. Table 10 evaluates the conditional probabilities between wrong code and wrong plan verification, as well as between correct code and correct plan verification. Typically, in LPW, accurate plan verifications significantly contribute to correct program solutions, whereas incorrect plan verifications inevitably result in errors. LPW generates program solutions based on plans and plan verifications. Therefore, any accurate descriptions or mistakes, including missed conditions, in the plan and plan verification are propagated to the code. When plan verifications are accurate, 96.1% of program solutions are correct in LPW. The remaining incorrect instances arise from unclear condition statements in plan verification that fail to account for hidden tests, leading to erroneous program solutions.
The results from Tables 9 and 10 highlight the impressive capability of LLMs in tackling text-to-code generation tasks when outputs are represented in natural language. Plan and plan verification generation accuracy is typically higher than code generation accuracy, underscoring the rationale behind LPW, which produces the high-quality program solution by leveraging plan and plan verification. It is worth exploring methods to help LLMs overcome the challenges of translating natural language solutions into programs, given the strict lexical, grammatical, and semantic constraints. Integrating alternative solution representations (Zelikman et al., 2023) alongside natural language representations could offer a promising approach.
Appendix F Refinement Consistency in LPW
@@
LPW allows multiple rounds of debugging to refine code based on error analysis, generated by comparing the code execution trace with plan verification on the failed visible test. LPW queries LLMs to generate refined code accompanied by a refinement explanation, detailing the modifications implemented to address the errors identified in the error analysis. For instance, Figures 14 and 7 illustrate two HumanEval problems where LPW successfully generates the correct program through refinements informed by the error analysis using the GPT-3.5 backbone. We note that in LPW, if the refined code is irrelevant to the error analysis, the entire debugging process degrades to a simple sampling approach, contradicting our original intent. As a result, we manually evaluate the debugging consistency among the generated error analysis (part (e)), the refined code (part (f)), and the refinement explanation (part (g)), as exampled in Figure 14. LPW demonstrates excellent consistency between error analysis and the refined code, where only one refined code deviates from the error analysis yet still produces the correct solution, across all problems solved through debugging. This result validates the effectiveness of the debugging steps in the code implementation phase for LPW, where the meaningful error analysis enables LLMs to produce the correct program with precise refinements.
Figure 12: Pass@1 accuracy of Baseline, LDB, and LPW across different difficulty levels, Easy, Medium, and Hard, on the LiveCode benchmark when using GPT-4o as the LLM backbone.
[…]
LDB 23.4
GPT-4o LPW (ours) 31.6
[…]
Table 11: The percentage of problems where LDB and LPW generated programs solve the visible tests but fail the hidden tests, out of total failed problems for each method in LiveCode, with GPT-4o as the backbone.
[…]
LDB 10.7
GPT-4o LPW (ours) 12.9
[…]
Table 12: The percentage of problems where LDB and LPW generated programs pass the visible tests but fail the hidden tests, out of a total of 140 problems in LiveCode, with GPT-4o as the backbone.
Appendix G Analysis of Unsolved Problems for LPW using GPT-4o
G.1 HumanEval
@@
LPW achieves 98.2% Pass@1 accuracy on HumanEval with the GPT-4o backbone, indicating only 3 unsolvable problems. We further investigate the reasons behind GPT-4o’s failures on the 91st, 132nd, and 145th problems as shown in Figures 15, 16, and 17, and attempt to generate the correct program solutions. The 91st problem fails since GPT-4o misinterprets the linguistic distinction between the word and the letter; the 132nd problem’s ambiguous description challenges GPT-4o; and the incomplete description of the 145th problem leads to failed plan verifications. LPW successfully generates correct program solutions for 2 out of 3 problems, achieving 99.4% Pass@1 accuracy, by involving an additional visible test to validate the intended solution for the 91st problem and providing a comprehensive problem description for the 145th problem.
G.1.1 The 91st Problem
@@ G.1.2 The 145th Problem
G.1.3 The 132nd Problem
Figure 16 illustrates the 132nd problem, where ambiguities in the problem description (part (a)) challenge GPT-4o. The problem description lacks clarity on “a valid subsequence of brackets” and fails to specify the meaning of “one bracket in the subsequence is nested”. We deduce the intended problem description by prompting GPT-4o with a canonical solution (part (d)). However, the LLM-generated description remains unclear and results in various erroneous programs. Furthermore, adding typically failed hidden tests to the visible tests (part (b)) is also ineffective in clarifying the correct logic. We acknowledge that a clearer description might help generate the correct program. However, some problems are inherently difficult to describe accurately in natural language without careful organization, posing challenges for LLMs.
@@ G.1.3 The 132nd Problem
Table 13: The percentage of problems where LDB and LPW generated programs solve the visible tests but fail the hidden tests, out of total failed problems for each method in APPS and CodeContests, with GPT-4o as the backbone.
@@
Table 14: The percentage of problems where LDB and LPW generated programs pass the visible tests but fail the hidden tests, out of a total of 139 problems in APPS and 150 problems in CodeContests, with GPT-4o as the backbone.
@@
Figure 13: Pass@1 accuracy of Baseline, LDB, and LPW across different difficulty levels, Introductory, Interview, and Competition, on the APPS benchmark when using GPT-4o as the LLM backbone.
G.2 LiveCode
LiveCode, a contamination-free dataset, serves as a reliable benchmark for evaluating code generation performance. It presents challenges for the advanced LLM GPT-4o, with all competing approaches showing performance limitations. We note that LPW consistently surpasses Baseline and LDB by approximately 15% and 5%, respectively, in Pass@1 accuracy and across different difficulty levels, as discussed below, emphasizing the reliability of LPW.
G.2.1 Performance Across Different Difficulty Levels
Figure 12 compares the Pass@1 accuracy of Baseline, LDB, and LPW across varying difficulty levels, Easy, Medium, and Hard on the LiveCode benchmark using GPT-4o. LPW achieves the highest Pass@1 accuracy across all levels, surpassing LDB by over 5% accuracy in each level. Compared to Baseline, LPW delivers over 15% higher accuracy at the Easy and Medium levels and 7.5% higher at the Hard level. Conversely, LDB performs similarly to Baseline at the Hard level, underscoring its limited refinement capability in addressing more complex tasks.
G.2.2 Failure on Hidden Tests
Tables 11 and 12 show the percentage of problems where the program solutions generated by LDB and LPW pass only the visible tests but fail the hidden tests out of the total number of failed problems and the total number of problems, respectively, on the LiveCode benchmark using GPT-4o. For LPW, 31.6% of failures occur when only the visible tests are solved, while for LDB, this percentage is 23.4%, as shown in Table 11. In Table 12, both LDB and LPW generate a similar proportion of solutions that pass only the visible tests. However, LPW tends to solve only the visible tests in 12.9% of problems, compared to 10.7% for LDB.
G.3 APPS and CodeContests
@@
APPS and CodeContests are unstructured benchmarks where visible tests are intermingled with the problem statements and function signatures are excluded. To align input data structure across benchmarks, we instruct GPT-4o to derive the optimal function signature and identify visible tests for each problem in APPS and CodeContests prior to conducting experiments. Example structured problems from APPS and CodeContests are illustrated in Figures 18 and 19. LPW demonstrates significant improvements on APPS and CodeContests, exceeding around 10% and 5% Pass@1 accuracy, respectively, compared to LDB with GPT-4o. However, in contrast to the performance on the HumanEval and MBPP benchmarks, where LPW achieves 98.2% and 84.8% Pass@1 accuracy, the 62.6% accuracy on APPS and 34.7% accuracy on CodeContests indicate that even for the advanced LLM GPT-4o, code generation remains challenging when addressing complicated programming problems, such as those encountered in collegiate programming competitions like IOI and ACM (Hendrycks et al., 2021).
G.3.1 Performance Across Different Difficulty Levels
@@
Figure 13 compares the Pass@1 accuracy of Baseline, LDB, and LPW across different difficulty levels, Introductory, Interview, and Competition, on the APPS benchmark using GPT-4o. LPW consistently dominates in Pass@1 accuracy across all difficulty levels. LPW shows strong performance on the relatively easier levels, i.e., Introductory and Interview, surpassing LDB by around 9% and 13% accuracy, respectively, and outperforming Baseline by over 20% accuracy. For the problems belonging to the most challenging level, Competition, LPW achieves 34.8% accuracy, compared to 28.3% for LDB and 17.4% for Baseline. However, all approaches experience a substantial decrease at the Competition level, emphasizing the necessity for further improvements.
G.3.2 Failure on Hidden Tests